OpenStudy (anonymous):

What are the asymptotes of the hyperbola given by the equation x^2/144 - y^2/4? Answer choices are; a. y=6x b. y=36x c. y=1/6x d. y=1/36x

jhonyy9 (jhonyy9):

so x/12 =y/2 2x=12y x=6y y=x/6 so c. is right

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

thanks (:

jhonyy9 (jhonyy9):

yw good luck bye